OpenStudy (anonymous):

HELP Find the markup on the following item. A necktie costing $30.00 per dozen and selling for $5 each (Be careful--find the cost of one tie before you subtract.)

OpenStudy (welshfella):

cost of one tie = 30/12 dollars markup = 5 - (30/12)

OpenStudy (anonymous):

Thank @welshfella

OpenStudy (welshfella):

yes work out the 30/12 first

OpenStudy (welshfella):

yw

OpenStudy (anonymous):

i will